OpenStudy (anonymous):

Solve. x2 – x = 0 x = 0 x = 1 and x = –1 x = 0 and x = –1 x = 0 and x = 1

OpenStudy (cwrw238):

first factor the LHS x(x - 1) = 0

OpenStudy (jhannybean):

\[\large x^2-x = 0\] factor out an x from this,since it your LCM.\[\large x(x-1)=0\] Now solve for both values. \[\large x=0\]\[\large x-1=0\]